OpenStudy (anonymous):

Help me please! In a triangle ABC, Angle B is 3 times Angle A. Angle C is 19 degrees less than 6 times Angle A. Find the size of the angles.

OpenStudy (anonymous):

B= 3 A C = 6A - 19 A + B + C = 180 A + 3A + 6A - 19 = 180 10A = 180 + 19 10 A = 199 => A = 19.9 degree B = 3A = 19.9 x 3 C = 6A - 19 = 6 (19.9) - 19

OpenStudy (anonymous):

\[A + B + C = 180\] \[A + 3*A + 6*A - 19 = 180\] \[10A = 199\] A = 19.9 B = 59.7 C = 100.4

OpenStudy (anonymous):

Thank you!